# Optimal Distributed Similarity Estimation for Unitary Channels

### v1

@@ III.1 Incoherent Access
The proof is provided in Appendix B.1. We remark that this incoherent algorithm naturally generalizes to estimating similarities between arbitrary channels. Concretely, we can estimate the inner product of two unknown channels’ Choi states [40], and then estimate the similarity via several inner product-based distance measures [24, 41]. Furthermore, this incoherent algorithm can also serve as a tool of unitarity estimation for an unknown quantum channel [38].
III.2 Coherent Access
We now turn to the coherent access and assume that two devices share only the randomness of state preparation. In various quantum learning scenarios, coherent access often provides substantial advantages compared to incoherent access [37, 38, 34, 39]. Here, we provide a coherent algorithm based on a symmetric collective measurement applied on quantum states in $\mathcal{H}^{\otimes T}$, labeled as $\mathcal{M}_{T}$ [42, 34]. $\mathcal{M}_{T}$ can be constructed as the following POVM:
$\mathcal{M}_{T}\mathrel{\mathop{\ordinarycolon}}=\left\{\frac{\kappa_{T}}{L}|{\phi_{j}}\rangle\!\langle{\phi_{j}}|^{\otimes T}\right\}_{j=1}^{L}\cup\{I-\Pi_{\rm sym}^{(d,T)}\},$ (5)
where $\{|{\phi_{j}}\rangle\}$ forms a state $(T+2)$-design, $\kappa_{T}\mathrel{\mathop{\ordinarycolon}}=\binom{d+T-1}{T}$, and $\Pi_{\rm sym}^{(d,T)}$ is the projector onto the symmetric subspace of $\mathcal{H}^{\otimes T}$. With this measurement, we provide the following algorithm:
  1. 1.
    Randomly generate pure states $|{\psi}\rangle^{\otimes T}$ on each quantum devices, where $|{\psi}\rangle$ from a state $4$-design ensemble;
  2. 2.
    Apply the unitary channels $\mathcal{U}^{\otimes T}$ and $\mathcal{V}^{\otimes T}$ on $|{\psi}\rangle^{\otimes T}$.
  3. 3.
    Measure $(\mathcal{U}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ and $(\mathcal{V}(|{\psi}\rangle\!\langle{\psi}|))^{\otimes T}$ with the POVM $\mathcal{M}_{T}$. Recode the measurement results on two devices as $|{\phi_{A}}\rangle$ and $|{\phi_{B}}\rangle$ and compute $\tilde{f}=|\langle\phi_{A}|\phi_{B}\rangle|^{2}$.
  4. 4.
    Obtain an unbiased estimator for $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$:
This algorithm is described in pseudocode in Algorithm 2. We summarize the query complexity of the coherent algorithm in the following theorem, proven in Appendix B.2.
Theorem 2.
The expectation and variance of estimator $\tilde{\chi}$ defined in Eq. (6) are given by
$\mathbb{E}\tilde{\chi}=\frac{1}{d^{2}}\operatorname{Tr}^{2}[U^{\dagger}V],\quad\mathbb{V}(\tilde{\chi})=\mathcal{O}\left(\frac{d}{m^{2}}+\frac{1}{m}+\frac{1}{d}\right).$ (7)
Surprisingly, although incoherent access is generally weaker than coherent access, the two algorithms have the same query complexity of completing DSEU. Let’s elaborate the differences between these two algorithms. First, unlike the incoherent algorithm, the coherent algorithm fundamentally requires that the unknown channels be unitary and the proof relies on pure output states; check out details in Appendix B.2. Second, the coherent algorithm does not reduce to the incoherent algorithm when $T=1$, as they have different classical data post-processing, caused by the different types of applied shared randomness. Specifically, the coherent algorithm requires sharing only the randomness of state preparation, whereas the incoherent algorithm additionally relies on shared measurement settings. This shared measurement randomness effectively boosts the performance of the incoherent algorithm, enabling it to achieve the $\mathcal{O}(\sqrt{d})$ query complexity.
IV Lower Bound
In this section, we provide the lower bounds for completing DSEU, showing that both two algorithms proposed in this work achieve optimal query complexity. To identify the lower bound of DSEU, we consider a related hypothesis testing task, which serves as a restricted version of DSEU. This is a task of distinguishing between the following two cases:
Two devices perform the same unitary $U$, which is a Haar random unitary;
Two devices independently perform two unitaries $U$ and $V$, which are both independent Haar random unitaries.
If a learning algorithm can estimate values of $\operatorname{Tr}^{2}[U^{\dagger}V]/d^{2}$, then, we can use this learning algorithm to complete the above hypothesis testing task. Hence, a lower bound for this task directly implies a lower bound for completing DSEU. Moreover, this lower bound also applies to estimating other commonly used distance measures between quantum channels. In the following, we analyze the query complexity required to solve this hypothesis testing task with incoherent and coherent accesses.
We employ Le Cam’s two-point method [43] to upper bound the success probability of solving the hypothesis testing problem. This method relates the success probability to the total variation distance (TVD) between the probability distributions of measurement results under the two hypotheses. Let $p_{1}(\ell)$ and $p_{2}(\ell)$ be the probabilities of obtaining measurement result $\ell$ after querying two unitary channels in the first and second cases. Their TVD is defined as
$\left\lVert p_{1}-p_{2}\right\rVert_{\rm TV}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{2}\sum_{\ell}\left|p_{1}(\ell)-p_{2}(\ell)\right|,$ (8)
To reliably distinguish the two hypotheses, this TVD must be at least constantly large. Now, we are ready to prove the lower bounds for two learning accesses.
IV.1 Incoherent Access
We first consider learning with incoherent access, and follow the framework of learning tree [37, 38, 44, 45, 46]. We show that for incoherent access, completing DSEU needs exponentially many queries, even learning with multi-round LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings. The result is shown in the following theorem.
Theorem 3.
For any algorithm with incoherent access that can be possible multi-round LOCC, ancilla-assisted, and adaptive, requires $\Omega(\sqrt{d})$ queries to complete DSEU.
This theorem shows that our incoherent algorithm in Section III.1 is optimal, yielding the query complexity of completing DSEU with incoherent coherent access is $\Theta(\sqrt{d})$. This scaling matches that of unitarity estimation [38], consistent with the fact that unitarity estimation is a special case of DSEU.
Proof sketch of Theorem 3.
Inspired by [38], we introduce an intermediate distribution $p_{\mathcal{D}}(\ell)$, which is the probability of obtaining measurement result $\ell$ when both channels are replaced by the completely depolarizing channel, defined as $\mathcal{D}(A)\mathrel{\mathop{\ordinarycolon}}=\operatorname{Tr}[A]I/d$. Then, with the triangle inequality, the target TVD is upper bounded by two TVDs: $\left\lVert p_{1}-p_{\mathcal{D}}\right\rVert_{\rm TV}$ and $\left\lVert p_{\mathcal{D}}-p_{2}\right\rVert_{\rm TV}$. Lastly, using techniques in [38], each of these terms is upper bounded by $\mathcal{O}(T^{2}/d)$, where $T$ is the number of queries made to each device. Therefore, completing DSEU with incoherent access requires query complexity $\Omega(\sqrt{d})$. The detailed proof is provided in Appendix C.1. ∎
IV.2 Coherent Access
We now turn to coherent access, where the learner may interleave multiple queries to the unknown channel with arbitrary quantum channels before the measurement. Perhaps surprisingly, we show that even under this significantly more powerful model, completing DSEU still requires exponentially many queries. This lower bound holds even when learning with multi-way LOCC, arbitrarily large ancillary systems, and adaptive SPAM settings. This result further highlights the fundamental hardness of DSEU. Our result is stated as follows.
Theorem 4.
For any algorithm with coherent access that can be possible multi-round LOCC, ancilla-assisted, and adaptive, requires $\Omega(\sqrt{d})$ queries to complete DSEU.
Thus our coherent algorithm in Section III.2 is also optimal, and the query complexity of completing DSEU with coherent access is also $\Theta(\sqrt{d})$. Consequently, coherent access offers no advantage for completing DSEU. This phenomenon is consistent with the result of distributed inner product estimation [15], which show that symmetric collective measurements do not reduce the complexity in distributed learning tasks. Taken together, these results highlight that shared randomness, rather than coherent access or collective measurements, is the key resource that enables efficient distributed learning.
Proof sketch of Theorem 4.
Assume we query $T$ times on each devices. Generally, analyzing the complexity of coherent access is subtle, as arbitrary quantum channels may appear between queries. To handle this difficulty, inspired by [47], we show that it suffices to work with the Choi operator of $\mathcal{U}^{\otimes T}$, defined as $J_{U}^{(T)}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{U}^{\otimes T}\otimes\mathcal{I})(|{\Phi}\rangle\!\langle{\Phi}|)$, where $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum|{ii}\rangle$ is the unnormalized maximally entangled states in $(\mathcal{H}^{\otimes T})^{\otimes 2}$. By formulating the problem in terms of Choi operators, all interleaving channels can be absorbed into an enlarged effective input system, reducing the analysis to distinguishing between the two averaged Choi operators
$\bar{J}_{1}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U}\;J_{U}^{(T)}\otimes\mathbb{E}_{V}\;J_{V}^{(T)},\quad\bar{J}_{2}^{(T)}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{U}\;J_{U}^{(T)}\otimes J_{U}^{(T)},$ (9)
which correspond to the two hypothesis testing cases.
Subsequently, as in the proof of lower bound for incoherent access, we introduce an intermediate probability distribution $p_{a}(\ell)$, defined as the probability of obtaining measurement result $\ell$ when both channels are replaced by the approximate $T$-design channel with Choi operator $J_{a}^{(T)}$ (see definition in [47], also Appendix Eq. (25)). We again analyze the TVDs: $\left\lVert p_{1}-p_{a}\right\rVert_{\rm TV}$ and $\left\lVert p_{a}-p_{2}\right\rVert_{\rm TV}$ separately. For the first term, we directly bound it using the relative error of the approximate $T$-design channel, and the corresponding result is $\mathcal{O}(T^{2}/d)$. For the second term, we relate it to the distance between the approximate $T$-design channel and a “measure-and-prepare” map. Using tools from quantum cloning theory [48],we show that this term is also upper bounded by $\mathcal{O}(T^{2}/d)$. Therefore, completing DSEU with coherent access requires query complexity $\Omega(\sqrt{d})$. The detailed proof is provided in Appendix C.2. ∎
V Independent Classical Shadow
@@
We now compare our algorithms with independently performing classical shadow for unitary channels [34], where no shared randomness is available. For completeness, we briefly outline how DSEU can be completed using independent classical shadow; the full procedure is given in Algorithm 3. For each device, we randomly select a state $|{\psi}\rangle$ in $\mathcal{H}$ from a state $4$-design ensemble, then apply $\mathcal{U}^{\otimes s}$ on $|{\psi}\rangle^{\otimes s}$, and finally perform the symmetric collective measurement $\mathcal{M}_{s}$. Repeating the above procedure $T$ times on each devices yields two independent sets of classical snapshots of $\mathcal{U}$ and $\mathcal{V}$, labeled as $\{X_{t}\}$ and $\{Y_{t}\}$. Lastly, the similarity can then be estimated via
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}[X_{i}^{\dagger}Y_{j}].$ (10)
Unlike our algorithms, this method uses no shared randomness, i.e., SPAM settings are chosen independently on the two devices. We provide the corresponding query complexity of this algorithm in the following theorem.
Theorem 5.
@@
The expectation and variance of estimator $\tilde{\gamma}$ defined in Eq. (10) are given by
@@
$\mathbb{E}\tilde{\gamma}=\frac{1}{d^{2}}\operatorname{Tr}^{2}[U^{\dagger}V],\quad\mathbb{V}(\tilde{\gamma})=\mathcal{O}\left(\frac{d^{2}}{T^{2}s^{2}}+\frac{1}{T}\right).$ (11)
@@
Consequently, the query complexity is $\mathcal{O}(\max\{1/\varepsilon^{2},d/\varepsilon\})$.
The proof is provided in Appendix D. This result shows that, regardless of how many queries are used to generate a single classical snapshot, completing DSEU with independent classical shadow requires $\mathcal{O}(d)$ queries. In contrast, both of our algorithms achieve query complexity $\Theta(\sqrt{d})$, revealing a square-root advantage that arises precisely from the shared randomness of SPAM settings.
VI Conclusion
In this work, we showed that the query complexity of completing DSEU is $\Theta(\sqrt{d})$, for both incoherent and coherent accesses. Firstly, we proposed two randomized measurement based estimation algorithms that effectively utilize shared randomness between the two quantum devices. Then, we proved that it is necessary and sufficient to take $\Theta(\sqrt{d})$ queries to complete DSEU. We also compared our algorithms with independent classical shadow, demonstrating a square-root advantage in query complexity. This advantage arises from the shared randomness employed in our algorithms, which allows the two devices to coordinate SPAM settings. The obtained results are summarized in Table 1.
Several questions remain open. For example, existing cross-platform verification protocols often assume that the two platforms implement the same operation, an assumption that may fail in realistic experimental settings due to device imperfections. Developing robust verification algorithms that relax this assumption is an important direction for future research.
Learning Access Upper Bound Lower Bound
Incoherent $\mathcal{O}(\sqrt{d})$ (Theorem 1) $\Omega(\sqrt{d})$ (Theorem 3)
@@ VI Conclusion
Coherent $\mathcal{O}(\sqrt{d})$ (Theorem 2) $\Omega(\sqrt{d})$ (Theorem 4)
Classical Shadow $\mathcal{O}(d)$ (Theorem 5)
Table 1: Our results on the query complexity of completing DSEU, where $d$ is the dimension of the unitary channels.
@@
While completing this manuscript, we became aware of a related work by Ananth et al. [49], which also studied the distinguishing problem constructed in our lower bound with coherent access. While both works identify the $\Omega(\sqrt{d})$ barrier, the proof techniques are distinct and complementary. Our approach relies on Le Cam’s two-point method and the connection to optimal quantum cloning theory [48], whereas Ananth et al. utilized linear programming. The diversity of proof techniques enriches the understanding of the fundamental hardness of this distinguishing problem.
@@ Acknowledgments.
This work was supported by the National Natural Science Foundation of China (Grant No. 62471126), the Jiangsu Frontier Technology Research and Development Plan (Grant No. BF2025066), the Fundamental Research Funds for the Central Universities (Grant No. 2242022k60001), the SEU Innovation Capability Enhancement Plan for Doctoral Students (Grant No. CXJH_SEU 24083), and the National Key R&D Program of China (Grant No. 2022YFF0712800).
@@ Appendix A Preliminaries
Let $\mathcal{H}$ be the Hilbert space of an $n$-qubit system, with dimension $d=2^{n}$. The set of Hermitian operators on $\mathcal{H}$ is denoted by $\mathcal{B}(\mathcal{H})$, the set of linear operators on $\mathcal{H}$ is denoted by $\mathcal{L}(\mathcal{H})$, the set of density matrices on $\mathcal{H}$ is denoted by $\mathcal{D}(\mathcal{H})$. For a pure state $|{\phi}\rangle\!\langle{\phi}|\in\mathcal{D}(\mathcal{H})$, we use the shorthand $\phi\mathrel{\mathop{\ordinarycolon}}=|{\phi}\rangle\!\langle{\phi}|$. For a unitary $U\in\mathcal{L}(\mathcal{H})$, let $\mathcal{U}(A)\mathrel{\mathop{\ordinarycolon}}=UAU^{\dagger}$ be the corresponding unitary channel. For a quantum channel $\mathcal{E}$, we define its Choi operator as [40]
$J_{\mathcal{E}}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{E}\otimes\mathcal{I})(\Phi),$ (12)
where $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum_{i}|{ii}\rangle$ is the unnormalized maximally entangled states. The Choi operator provides the useful identity
$\operatorname{Tr}\left[O\mathcal{E}(\rho)\right]=\operatorname{Tr}\left[\left(O\otimes\rho^{T}\right)J_{\mathcal{E}}\right],$ (13)
which we will frequently employ throughout this work.

### v2

@@ Abstract
Combined with Theorem 3, the query complexity of DSEC is exactly $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$, as this incoherent algorithm matches the lower bound and applies to general channels.
This incoherent algorithm enjoys several notable advantages. First, it is ancilla-free and non-adaptive, making it immediately implementable on near-term devices. Second, it can also serve as a tool for unitarity estimation [41], since it works without the unitary assumption. Third, despite incoherent access being strictly weaker than coherent access in general, both algorithms achieve the same query complexity for DSEC. This surprising phenomenon parallels the result for DIPE [15], which shows that symmetric collective measurements do not reduce complexity in distributed learning.
The key to these advantages lies in the different types of shared randomness employed. The coherent algorithm shares only the randomness for state preparation, whereas the incoherent algorithm additionally shares measurement settings. This shared measurement randomness effectively boosts the performance of the incoherent algorithm, enabling it to achieve the optimal query complexity for general channels. This finding highlights that shared randomness, rather than coherent access or collective measurements, is the key resource for efficient distributed learning. On the other hand, the coherent algorithm has its own advantage: less shared randomness translates to less communication overhead. Thus, there exists a trade-off between classical communication resources and the power of the learning algorithm.
Classical Shadow.— To quantify the advantage of shared randomness, we compare our algorithms with independent classical shadow for channels [35, 36, 37], which use no shared randomness between devices. We adopt the approach of [37] as our baseline, since it achieves near-optimal performance for unitary channels and naturally extends to incoherent access for general channels.
@@
The independent classical shadow approach for DSEC proceeds as follows (see Algorithm 3 for details). Suppose $\mathcal{E}(\cdot)=U(\cdot)U^{\dagger}$ and $\mathcal{F}(\cdot)=V(\cdot)V^{\dagger}$ are unitary channels. On each device independently, we randomly select a state $|{\psi}\rangle$ from a state $4$-design ensemble, apply $U^{\otimes s}$ (or $V^{\otimes s}$) to $|{\psi}\rangle^{\otimes s}$, and perform the symmetric collective measurement $\mathcal{M}_{s}$. Repeating this procedure $T$ times on each device yields two independent sets of classical snapshots, $\{X_{t}\}$ and $\{Y_{t}\}$. The inner product is then estimated via
$\tilde{\gamma}=\frac{1}{T^{2}d^{2}}\sum_{i,j=1}^{T}\operatorname{Tr}[X_{i}Y_{j}].$ (10)
The key difference from our algorithms is that no shared randomness is used; SPAM settings are chosen independently on each device. The following proposition, proven in Appendix D, establishes the query complexity.
Proposition 6.
@@
For unitary channels $\mathcal{E}$ and $\mathcal{F}$, the expectation and variance of estimator $\tilde{\gamma}$ defined in Eq. (10) are given by
@@
$\mathbb{E}\tilde{\gamma}=\frac{1}{d^{2}}\operatorname{Tr}[J_{\mathcal{E}}J_{\mathcal{F}}],\quad\mathbb{V}(\tilde{\gamma})=\mathcal{O}\left(\frac{d^{2}}{T^{2}s^{2}}+\frac{1}{T}\right).$ (11)
@@
Consequently, the query complexity is $\mathcal{O}(\max\{d/\varepsilon,1/\varepsilon^{2}\})$. For $s=1$, the result extends to general channels.
This result reveals that independent classical shadow requires $\mathcal{O}(\max\{d/\varepsilon,1/\varepsilon^{2}\})$ queries for DSEC, regardless of the number of queries per snapshot. Notably, the lower bound in Theorem 2 also applies to independent classical shadow based algorithms, as they fall within the same learning access. In contrast, our algorithms with shared randomness achieve the optimal query complexity of $\mathcal{O}(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$, demonstrating a quadratic improvement in the dimension-dependent term. This square-root advantage directly arises from the shared randomness in SPAM settings, which enables the two devices to coordinate their measurements.
Access Model Lower Bound Algorithm Query Complexity Applicable Channel
@@
Coherent $\Omega(\sqrt{d})$ (Theorem 3) Independent classical shadow $\mathcal{O}(d)$ (Proposition 6) Unitary
RM with shared randomness $\mathcal{O}(\sqrt{d})$ (Theorem 4) Unitary
Incoherent Independent classical shadow $\mathcal{O}(d)$ (Proposition 6) General
RM with shared randomness $\mathcal{O}(\sqrt{d})$ (Theorem 5) General
Table 1: Summary of our query complexity results for distributed similarity estimation of quantum channels (DSEC), where $d$ is the channel dimension and RM is short for randomized measurements.
Conclusion.— In this work, we characterized the optimal query complexity of distributed similarity estimation of quantum channels (DSEC). We proved that $\Theta(\max\{\sqrt{d}/\varepsilon,1/\varepsilon^{2}\})$ queries are both necessary and sufficient for this task. The lower bound holds even in the strongest setting (multi-round LOCC, coherent access, and adaptive strategies), while the upper bound is achieved in the weakest setting (incoherent access, ancilla-free, and non-adaptive). For the upper bound, we proposed two randomized measurement-based algorithms that effectively utilize shared randomness between devices. Both achieve optimal query complexity, with the incoherent algorithm applicable to general channels. We also compared our algorithms with independent classical shadow, demonstrating a square-root advantage in query complexity. This advantage arises from the shared randomness employed in our algorithms, which allows the two devices to coordinate SPAM settings. The obtained results are summarized in Table 1.
Several directions remain open for future research. First, existing cross-platform verification protocols typically assume identical operations on both platforms, which may fail in practice; developing robust verification algorithms that relax this assumption is an important goal. Second, since classical shadow methods utilize information about sampled unitaries, investigating whether this information can be better exploited to reduce query complexity is an intriguing question.
@@
While completing this manuscript, we became aware of a related work by Ananth et al. [55], which also studied the lower bound for completing Problem 1 with coherent access. While both works identify the $\Omega(\sqrt{d})$ barrier, the proof techniques are distinct and complementary. Our approach relies on Le Cam’s two-point method and connections to optimal quantum cloning [51], whereas Ananth et al. utilize linear programming. This diversity of techniques enriches our understanding of the fundamental hardness of this problem. In addition, we establish the $\varepsilon$-dependent lower bound.
@@ Acknowledgments.
This work was supported by the National Natural Science Foundation of China (Grant Nos. 62471126 and 12504584), the Jiangsu Frontier Technology Research and Development Plan (Grant No. BF2025066), the Fundamental Research Funds for the Central Universities (Grant No. 2242022k60001), the SEU Innovation Capability Enhancement Plan for Doctoral Students (Grant No. CXJH_SEU 24083), and the National Key R&D Program of China (Grant No. 2022YFF0712800).
@@ Appendix A Preliminaries
In this section, we first set some notation used in this work. Then, we introduce two central tools utilized in this work: the Haar measure and approximate unitary designs. Both play fundamental roles in quantum information theory [56, 50], and they are essential for establishing the query complexities of our algorithms as well as the lower bounds for completing DSEC.
A.1 Notation
Let $\mathcal{H}$ be the Hilbert space of a $d$-dimensional system. The set of Hermitian operators on $\mathcal{H}$ is denoted by $\mathcal{B}(\mathcal{H})$, and the set of density matrices on $\mathcal{H}$ is denoted by $\mathcal{D}(\mathcal{H})$. For a quantum channel $\mathcal{E}$, we define its Choi operator as [39]
$J_{\mathcal{E}}\mathrel{\mathop{\ordinarycolon}}=(\mathcal{E}\otimes\mathcal{I})(|{\Phi}\rangle\!\langle{\Phi}|),$ (12)
where $|{\Phi}\rangle\mathrel{\mathop{\ordinarycolon}}=\sum_{i}|{ii}\rangle$ is the unnormalized maximally entangled state. The Choi operator provides the useful identity
$\operatorname{Tr}\left[O\mathcal{E}(\rho)\right]=\operatorname{Tr}\left[\left(O\otimes\rho^{T}\right)J_{\mathcal{E}}\right].$ (13)
Additionally, we can also use a set of Kraus operators to describe a quantum channel. A quantum channel $\mathcal{E}$ can be described by a set of Kraus operators $\{E_{i}\}_{i}$ [57], and we have
$\mathcal{E}(\rho)=\sum_{i}E_{i}\rho E_{i}^{\dagger}.$ (14)
For two quantum channels $\mathcal{E}$ and $\mathcal{F}$, with Kraus operators $\{E_{i}\}_{i}$ and $\{F_{j}\}_{j}$, the inner product of their Choi operators can be represented as
$\operatorname{Tr}\left[J_{\mathcal{E}}J_{\mathcal{F}}\right]=\sum_{i,j}\operatorname{Tr}^{2}\left[E_{i}^{\dagger}F_{j}\right].$ (15)